# Can Finetuing LLMs on Small Human Samples Increase Heterogeneity, Alignment, and Belief-Action Coherence?

### v1

@@ 1 Introduction
With the rapid advancement of reasoning and language capabilities in large language models (LLMs), researchers have increasingly explored their use as human surrogates to simulate survey and experimental responses. One of the earliest works in this direction is by argyle2023out, who used GPT-3 to model public opinion patterns in political science. Since then, the use of LLMs as proxies for human participants has expanded across a range of disciplines, including marketing, economics, and psychology (e.g., brand2023using; aher2023using ), as well as in natural language processing (NLP) tasks traditionally performed by humans, such as text quality evaluation and data annotation (Chiang2023CanLL; he2024annollm). In parallel, another line of work has focused on finetuning open-source LLMs with human experimental data to better align model predictions with real behavioral distributions. For example, Centaur (binz2025foundation) is an LLM finetuned on a large-scale psychology dataset to more accurately mimic differences in human behavior, while SubPOP (suh2025language) leverages real survey data to model subpopulation opinion distributions. Together, these developments have spurred growing interest in whether LLMs can serve not only as reasoning systems but also as credible simulators of human heterogeneity and social behavior.
@@
The aim of this study is not to enter the discourse of whether off-the-shelf or pre-finetuned LLMs can fully substitute for human respondents, but rather to examine whether finetuning on a small subset of human data, such as that obtainable from a pilot study, can mitigate the limitations documented in prior work, such as the lack of heterogeneity (Gao), alignment (wang2025large), value-action coherence (shen2025mind) , and regression coefficient recovery. We base our simulations on the behavioral experiment of Hunt, which investigates how information disclosure shapes attackers’ judgments and decisions in a security scenario. Our findings show that LLMs finetuned on small fractions of the human dataset can substantially improve heterogeneity, distributional alignment, and value–action coherence relative to the base LLM. However, despite these improvements, none of the finetuned models reliably reproduces the original regression coefficients or underlying hypothesis tests, indicating that replacing human participants with LLMs remains risky and may introduce non-negligible bias into conclusions. We posit that the most promising use of LLMs in this context is in-silico prototyping for experimental design, helping researchers explore which treatments are likely to yield detectable effects, refine survey instruments, and estimate sample sizes, rather than serving as a direct replacement for human subjects (anthisposition; dillion2023can; binz2025foundation). In sum, we contribute to the literature by showing that fine-tuning on a small sample of human data (30 observations) can substantially mitigate several issues identified in prior work, including the lack of heterogeneity, distributional misalignment, and the value–action gap.
@@ 2.2 Attempts to Enhance the Behavioral Realism of LLMs
3 Study Setting
@@
This study builds on an large behavioral experiment by Hunt, which examined the effects of defensive information disclosure on attackers’ judgments and decisions in a security setting. Specifically, the experiments studied how information disclosed about the deployment of new security technology impacts attackers’ beliefs regarding where the technology is deployed and their final attack decisions (i.e., whether to attack, and if so, where). We selected this experiment for use in the current research for three main reasons. First, the design of the experiment, which elicited both beliefs and decisions from the participants, enables us to assess the coherence between LLMs’ stated beliefs and their subsequent actions. Second, we obtained the original data directly from the authors of the experiment, with confirmation that the data had not yet been publicly released or fed into any LLM prior to the analysis presented herein. Ensuring that the model had not previously seen the dataset is critical, as prior work demonstrates that LLMs can memorize training data, including personally identifiable information, and reproduce it verbatim with only one related document in the training corpus (zhou2024quantifying; carlini2021extracting; carlini2022quantifying). Such data contamination undermines the validity of LLM evaluations by producing spuriously high or unreliable performance estimates (sainz2023nlp; li2024open; ranaldi2024investigating). Third, the experiments in Hunt focused on information disclosure in a high-stakes, strategic environment (i.e., security and defense), a domain that has not been studied in the context of LLMs and behavioral realism.
@@
The sample we analyze consists of behavioral data from 929 U.S. adults  [^1] The original dataset in Hunt had 975 responses, but we exclude data points with missing demographic information for both finetuning and inference. who played the role of attackers deciding whether to attack either of two targets that were defended by a centralized defender. In the experiment, the defender released one of four disclosures related to the deployment of new security measures to protect her assets at the targets. The four different disclosures participants could have received are as follows: the defender stated the security was deployed (i) only at the first target, (ii) only at the second target, or (iii) at both targets, or (iv) the defender included no target-level information in the disclosure and just stated new technology was deployed. Participants (i.e., attackers) were also informed of the defender’s relative valuations of both targets, where the differences between her valuations at each target could be large (she values one target much more than the other) or small (she values one target only a bit more than the other). Appendix B outlines the manipulations in more detail. This 2 x 4 design created a total of eight treatment groups (i.e., two levels of the valuation factor, and four levels of the disclosure factor). After receiving the experimental manipulations, participants then reported their beliefs regarding where the technology was deployed – expressed as probabilities that the new security measures were deployed at Target 1, Target 2, or both (we refer to the probabilities (beliefs) as $\phi_{1}$, $\phi_{2}$, and $\phi_{B}$, respectively, throughout the remainder of this paper) – and then made a final decision on whether to attack Target 1, attack Target 2, or not attack. Participants were incentivized with base pay for completing the study and potential bonus payments based on the outcomes of their attacks. The full experimental script can be found in Appendix A.
4 Simulation
@@
All simulations were conducted using GPT-4.1 (achiam2023gpt) with the default temperature setting of 1.0. The finetuned models use the same base architecture and temperature configuration to ensure comparability. In the following subsection, we examine how finetuning affects heterogeneity, distributional alignment, and the belief–action gap, as well as the accuracy of regression coefficient estimates relative to the regression analysis performed on the human data.
@@ 4.1 Heterogeneity
To establish a baseline, we first ran simulations using an unfinetuned GPT model, since most prior studies evaluate only base LLMs without task-specific adaptation. In addition to the experimental questions, we incorporated participant demographics as role prompts to simulate more realistic responses.
4.1.1 Analysis of $\phi_{1}$
Figure 1 shows the distribution of beliefs regarding the likelihood that the defender’s security measures were deployed at Target 1 ($\phi_{1}$) and both targets ($\phi_{B}$). The analysis of Target 2’s beliefs ($\phi_{2}$) is presented in Appendix C and is very similar to what we uncover in this section. In analyzing $\phi_{1}$ in Figure 1, the unfinetuned model ("Unfinetuned baseline") exhibits a highly concentrated distribution, with most responses clustered in the 30–40 % and 40–50 % bins. In contrast, human responses follow a more bell-shaped pattern, with probability mass spread across multiple bins. This supports prior findings that base LLMs lack response heterogeneity. In an attempt to mitigate this issue, we finetune the model on a small fraction of the human data to examine whether heterogeneity would increase. Specifically, we finetune the model with just 25% of the data from a single treatment group (the V1A1 treatment group; see Appendix B), equating to 30 of the 929 observations.
@@ 4.1.1 Analysis of $\phi_{1}$
When we finetune the model on 25 % of the first treatment group (titled "Random (1a1b)" in Figure 1), the distribution becomes substantially more dispersed, showing probability mass across all bins and modest peaks around 50–70 %, closely resembling human patterns. Expanding finetuning to include randomly sampled data from four treatment groups ("Random (4 Groups)") and all eight groups ("Random (All Groups)") further improves diversity, with more formal alignment analysis discussed in the following subsection. In addition to random sampling, we also implemented balanced finetuning, drawing equal proportions [^2] We use a round-robin sampling strategy in which we randomly draw one sample from each ethnicity group in sequence until reaching 25% of the dataset. For ethnicity categories with smaller sample sizes (e.g., Other), we include all available observations and then continue the round-robin procedure with the remaining groups to reach the target proportion. of responses from each ethnic subgroup (25% of the samples) to mitigate the systematic misalignment for minority subgroups issue noted in prior research. These balanced variants, "Balanced (1a1b)," "Balanced (4 Groups)," and "Balanced (All Groups)," also show similarly enhanced diversity and alignment. This indicates that even a small number of finetuned samples can drastically improve diversity and mitigate homogeneity issues.
4.1.2 Analysis of $\phi_{B}$
As shown in Figure 1 (right panel), the human distribution for the belief that technology is deployed at both targets ($\phi_{B}$) is highly right-skewed, with most probability mass clustered near 0–10%. The finetuned models exhibit similar patterns as the human data and capture heterogeneity, but similar to the analysis of $\phi_{1}$, the unfinetuned baseline suffers from homogeneity, with the probability mass across all simulation in the 30–50% bins. This suggests that the base (unfinetuned) LLM does not place proper weight on the contextual variables provided in the study (e.g., valuations, disclosure type) and instead relies on a strong, low-variance prior across the three belief options. To further investigate this, instead of analyzing only one belief at a time, we also consider the entire belief structure – that is, the combinations of $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$ – given that each participant in the experiment, and each LLM simulation, recorded these three beliefs as part of a single response. When analyzing the belief structures elicited by humans and LLM, we find that the unfinetuned LLM suffers from extreme homogeneity. As shown in Table 1, the unfinetuned model produces only 19 unique belief structure (i.e., combinations of $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$) across all 929 simulations, compared to 340 in the human data. All finetuned models produce at least 200 unique combinations, indicating that even limited finetuning substantially increases the diversity of belief patterns expressed by the model.
@@ 5 Conclusion
This study examines whether finetuning on small samples of human participant data can mitigate well-documented limitations of using base LLMs to simulate survey responses (Gao; wang2025large; shen2025mind; bisbee2024synthetic), using a behavioral experiment on information disclosure from Hunt. Our analysis focuses on four key constructs: heterogeneity, distributional alignment, the value–action gap, and regression coefficient recovery. We find that finetuning on relatively small subsets of data substantially increases heterogeneity, improves alignment, and enhances belief–action coherence compared to the unfinetuned model, as demonstrated through both visual inspection and formal divergence metrics. Notably, heterogeneity can be substantially improved using as little as 25% of a single treatment group, approximately 30 observations. Alignment tends to strengthen as the number of treatment groups used for finetuning increases. Unexpectedly, however, sampling equal proportions from ethnicity subgroups does not consistently enhance alignment for minority groups relative to random sampling, and misalignment remains lowest for White participants even under balanced sampling. This pattern aligns with prior work showing that LLMs are more closely aligned with the WEIRD population (liu2022quantifying). In contrast, when examining the value–action gap, balanced sampling performs consistently better than random sampling, suggesting that belief–action coherence may be easier to steer than subgroup-level distributional alignment, which may require more advanced finetuning strategies.
@@ 6 Discussion
Several mechanisms may explain the misalignments between LLM-simulated and human responses. wang2025large argue that LLMs often engage in misportraying, rather than representing the authentic behaviors of a given demographic group, models reproduce what out-group members believe those in-group members would do. Since demographic identity is rarely encoded directly in the text on which LLMs are trained, explicit mentions of identity may instead express stereotypes by out-group authors. liu2024large offer a related explanation. They study LLM performance on datasets involving risky decision-making and find that LLMs align more closely with people’s expectations of how others behave than with how individuals actually behave. Since people tend to assume that others are more rational than themselves, LLMs may encode these perceptions rather than genuine human behavior. These findings suggest that LLM decisions may be shaped less by representations of actual in-group behavior and more by culturally shared beliefs about what “typical” behavior should look like. Additional factors may also contribute to misalignment. The training data themselves may contain survey responses affected by social desirability bias (crowne1960new), meaning that models learn socially acceptable patterns of responding rather than authentic responses. More directly, reinforcement learning from human feedback (RLHF) (ouyang2022training) optimizes models to behave in ways that humans prefer LLMs to behave (dahlgren2025helpful), not necessarily in ways that mirror empirical human data.

### v2

@@ 1 Introduction
With the rapid advancement of reasoning and language capabilities in large language models (LLMs), researchers have increasingly explored their use as human surrogates to simulate survey and experimental responses. One of the earliest works in this direction is by argyle2023out, who used GPT-3 to model public opinion patterns in political science. Since then, the use of LLMs as proxies for human participants has expanded across a range of disciplines, including marketing, economics, and psychology (e.g. brand2023using; aher2023using), as well as in natural language processing (NLP) tasks traditionally performed by humans, such as text quality evaluation and data annotation (Chiang2023CanLL; he2024annollm). In parallel, another line of work has focused on finetuning open-source LLMs with human experimental data to better align model predictions with real behavioral distributions. For example, Centaur (binz2025foundation) is an LLM finetuned on a large-scale psychology dataset to more accurately mimic differences in human behavior, while SubPOP (suh2025language) leverages real survey data to model subpopulation opinion distributions. Together, these developments have spurred growing interest in whether LLMs can serve not only as reasoning systems but also as credible simulators of human heterogeneity and social behavior.
@@
The aim of this study is not to enter the discourse of whether off-the-shelf or pre-finetuned LLMs can fully substitute for human respondents, but rather to examine whether finetuning on a small subset of human data, such as that obtainable from a pilot study, can mitigate the limitations documented in prior work. For reasons detailed below, we base our simulations on the behavioral experiment of Hunt, which investigates how information disclosure shapes attackers’ judgments and decisions in a security scenario. Our findings show that LLMs finetuned on small fractions of the human dataset can substantially improve heterogeneity, distributional alignment, and value–action coherence relative to the base LLM. However, despite these improvements, none of the finetuned models reliably reproduce the original regression coefficients or underlying hypothesis tests, indicating that replacing human participants with LLMs remains risky and may introduce non-negligible bias into conclusions. We posit that the most promising use of LLMs in this context is in-silico prototyping for experimental design, helping researchers explore which treatments are likely to yield detectable effects, refine survey instruments, and estimate sample sizes, rather than serving as a direct replacement for human subjects (anthisposition; dillion2023can; binz2025foundation). In sum, we contribute to the literature by showing that fine-tuning on a small sample of human data (30 observations) can substantially mitigate several issues identified in prior work, including the lack of heterogeneity, distributional misalignment, and the value–action gap. Both the original and simulated data will be released for future research upon acceptance.
@@ 2.2 Attempts to Enhance the Behavioral Realism of LLMs
2.3 Gaps in Existing Work
Despite substantial work documenting the limitations of LLMs as simulated survey respondents, very little research has explored whether these issues can be mitigated through targeted finetuning on small samples of human data. Existing studies typically evaluate base models or rely on prompt engineering, or persona conditioning, but do not examine whether a small, realistic amount of human data such as that obtainable from a pilot study can meaningfully improve heterogeneity, alignment, and behavioral coherence. This gap leaves open the critical question of whether modest finetuning can address the flaws identified in earlier research, or whether these limitations are inherent to current LLM architectures.
3 Data
@@
This study builds on a large behavioral experiment by Hunt, which examined the effects of information disclosure on attackers’ judgments and decisions in a security setting. Specifically, the experiments studied how information disclosed about the deployment of new security technology impacts attackers’ beliefs regarding where the technology is deployed and their final attack decisions (i.e., whether to attack, and if so, where). We selected this experiment for use in the current research for three main reasons. First, the design of the experiment, which elicited both beliefs and decisions from the participants, enables us to assess the coherence between LLMs’ stated beliefs and their subsequent actions. Second, we obtained the original data directly from the authors of the experiment, with confirmation that the data had not yet been publicly released or fed into any LLM prior to the analysis presented herein. Ensuring that the model had not previously seen the dataset is critical, as prior work demonstrates that LLMs can memorize training data, including personally identifiable information, and reproduce it verbatim with only one related document in the training corpus (zhou2024quantifying; carlini2021extracting; carlini2022quantifying). Such data contamination undermines the validity of LLM evaluations by producing spuriously high or unreliable performance estimates (sainz2023nlp; li2024open; ranaldi2024investigating). Third, the experiments in Hunt focused on information disclosure in a high-stakes, strategic environment (i.e., security and defense), a domain that has not been studied in the context of LLMs and behavioral realism.
@@
The sample we analyze consists of behavioral data from 929 U.S. adults  [^1] The original dataset in Hunt had 975 responses, but we exclude data points with missing demographic information for both finetuning and inference. who played the role of attackers deciding whether to attack either of two targets that were defended by a centralized defender. In the experiment, the defender released one of four disclosures related to the deployment of new security measures to protect her assets at the targets. Participants (i.e., attackers) were also informed of the defender’s relative valuations of both targets, where the differences between her valuations of each target could be large or small. Appendix B outlines the manipulations in more detail. This 2 x 4 design created a total of eight treatment groups (i.e., two levels of the valuation factor, and four levels of the disclosure factor). After receiving the experimental manipulations, participants then reported their beliefs regarding where the technology was deployed – expressed as probabilities that the new security measures were deployed at Target 1, Target 2, or both, and then made a final decision on whether to attack Target 1, attack Target 2, or not attack. Participants were incentivized with base pay for completing the study and potential bonus payments based on the outcomes of their attacks. The full experimental script can be found in Appendix A.
4 Methods
Our goal is to explore how well various LLM-based approaches can predict various results derived from this experimental dataset, and in particular how an approach fine-tuned on a subset of this data (e.g. that would come from a pilot study) can improve these predictions. We consider three different approaches, described below. All simulations were conducted using GPT-4.1 (achiam2023gpt) with the default temperature setting of 1.0. The finetuned models use the same base architecture and temperature configuration to ensure comparability.
Finetuning on experimental data
We used two sampling strategies and three dataset sizes, resulting in six models in total. The first strategy uses random sampling, and the second strategy, which we refer to as balanced sampling, draws equal proportions  [^2] We use a round-robin sampling strategy in which we randomly draw one sample from each ethnicity group in sequence until reaching 25% of the dataset. For ethnicity categories with smaller sample sizes (e.g., Other), we include all available observations and then continue the round-robin procedure with the remaining groups to reach the target proportion. of responses from each ethnic subgroup in order to address the systematic misalignment for minority groups noted in prior research. The three dataset sizes correspond to the number of treatment groups included in finetuning. This design allows us to test whether finetuning on only a subset of treatment groups enables the model to generalize to the remaining groups. Specifically, we sample 25 percent of the data from the first treatment group (1a1b), from the first four treatment groups, and from all eight treatment groups. For the multi group settings, we draw 25 percent from each included group. We name each model using its sampling strategy and dataset size. For example, the model finetuned on 25 percent of the first treatment group using random sampling is labeled Random (1a1b), with parallel naming for the remaining random and balanced models.
@@ 4 Methods
As outcomes, we examine how finetuning affects heterogeneity, distributional alignment, and the belief–action gap, as well as the accuracy of regression coefficient estimates relative to the regression analysis performed on the human data.
5 Results
In the following subsections, we examine how finetuning affects heterogeneity, distributional alignment, and the belief–action gap, as well as the accuracy of regression coefficient estimates relative to the regression analysis performed on the human data.
@@ 5.1 Heterogeneity
We find that even a small number of finetuned samples can drastically improve diversity and mitigate homogeneity issues relative to the unfinetuned baseline. Figure 1 shows the distribution of beliefs regarding the likelihood that the defender’s security measures were deployed at Target 1 and both Targets. The analysis of Target 2’s beliefs and the results from all models are presented in Appendix C, and support findings discussed here. In analyzing Target 1 (left panel) in Figure 1, the unfinetuned baseline exhibits a highly concentrated distribution, with most responses clustered in the 30–40 % and 40–50 % bins. In contrast, human responses follow a more bell-shaped pattern, with probability mass spread across multiple bins. This supports prior findings that base LLMs lack response heterogeneity. In an attempt to mitigate this issue, we finetune the model on a small fraction of the human data to examine whether heterogeneity would increase. Specifically, we finetune the model with just 25 % of the data from a single treatment group (the 1a1b treatment group; see Appendix B), equating to 30 of the 929 observations. When we finetune the model on 25 % of the first treatment group, the distribution becomes substantially more dispersed, showing probability mass across all bins and modest peaks around 50–70 %, closely resembling human patterns.
@@
As shown in Figure 1 (right panel), the human distribution for the belief that technology is deployed at both targets is highly right-skewed, with most probability mass clustered near 0–10%. The finetuned models exhibit similar patterns as the human data and capture heterogeneity, but similar to the analysis of Target 1, the unfinetuned baseline suffers from homogeneity, with the probability mass across all simulation in the 30–50% bins. This suggests that the unfinetuned baseline does not place proper weight on the contextual variables provided in the study (e.g., valuations, disclosure type) and instead relies on a strong, low-variance prior across the three belief options. To further investigate this, instead of analyzing only one belief at a time, we also consider the entire belief structure, given that each participant in the experiment, and each LLM simulation, recorded these three beliefs as part of a single response. When analyzing the belief structures elicited by humans and LLM, we find that the unfinetuned LLM suffers from extreme homogeneity. As shown in Table 1, the unfinetuned model produces only 19 unique belief structure across all 929 simulations, compared to 340 in the human data. All finetuned models produce at least 200 unique combinations, indicating that even limited finetuning substantially increases the diversity of belief patterns expressed by the model.
@@ 6 Conclusion
This study examines whether finetuning on small samples of human participant data can mitigate well-documented limitations of using base LLMs to simulate survey responses (Gao; wang2025large; shen2025mind; bisbee2024synthetic), using a behavioral experiment. Our analysis focuses on four key constructs: heterogeneity, distributional alignment, the value–action gap, and regression coefficient recovery. We find that finetuning on relatively small subsets of data substantially increases heterogeneity, improves alignment, and enhances belief–action coherence compared to the unfinetuned model, as demonstrated through both visual inspection and formal divergence metrics. Notably, heterogeneity can be substantially improved using as little as 25% of a single treatment group, approximately 30 observations. Alignment tends to strengthen as the number of treatment groups used for finetuning increases. Unexpectedly, however, sampling equal proportions from ethnicity subgroups does not consistently enhance alignment for minority groups relative to random sampling, and misalignment remains lowest for White participants even under balanced sampling. While unexpected, this pattern aligns with prior work showing that LLMs are more closely aligned with the WEIRD population (liu2022quantifying). In contrast, when examining the value–action gap, balanced sampling performs consistently better than random sampling, suggesting that belief–action coherence may be easier to steer than subgroup-level distributional alignment, which may require more advanced finetuning strategies.
@@ 7 Discussion
Several mechanisms may explain the misalignments between LLM-simulated and human responses. wang2025large argue that LLMs often engage in misportraying, rather than representing the authentic behaviors of a given demographic group, models reproduce what out-group members believe those in-group members would do. Since demographic identity is rarely encoded directly in the text on which LLMs are trained, explicit mentions of identity may instead express stereotypes by out-group authors. liu2024large offer a related explanation, and it aligns with the sociolingustic theoretical model introduced by [https://arxiv.org/abs/2502.12057]. They study LLM performance on datasets involving risky decision-making and find that LLMs align more closely with people’s expectations of how others behave than with how individuals actually behave. Since people tend to assume that others are more rational than themselves, LLMs may encode these perceptions rather than genuine human behavior. These findings suggest that LLM decisions may be shaped less by representations of actual in-group behavior and more by culturally shared beliefs about what “typical” behavior should look like. Additional factors may also contribute to misalignment. The training data themselves may contain survey responses affected by social desirability bias (crowne1960new), meaning that models learn socially acceptable patterns of responding rather than authentic responses. More directly, reinforcement learning from human feedback (RLHF) (ouyang2022training) optimizes models to behave in ways that humans prefer LLMs to behave (dahlgren2025helpful), not necessarily in ways that mirror empirical human data.
@@ Appendix C Belief
As shown in Figure 4, expanding finetuning to include randomly sampled data from four treatment groups ("Random (4 Groups)") and all eight groups ("Random (All Groups)") further improves diversity. The balanced variants, "Balanced (1a1b)," "Balanced (4 Groups)," and "Balanced (All Groups)," also show similarly enhanced diversity and alignment.